OpenStudy (anonymous):

What is the value of 2|6 + p| - 3p3 when p = -2?

OpenStudy (anonymous):

value = 2(6 - 2) - 3(-2^3) = 8 +24 = 32

OpenStudy (anonymous):

thanks :D